A force of 0.037 N is required to move a charge of 35.9 µC a distance of 24 cm in an electric field. What is the size of the potential difference between the two points?
1
Expert's answer
2020-02-20T09:58:46-0500

We write an expression for the work on the movement of a charge in an electric field

"A=q(\phi_2-\phi_1)=q \cdot\Delta\phi"

The same work can be expressed as

"A=F\cdot S"

then

"F\cdot S=q \cdot\Delta\phi"

from here

"\Delta\phi=\frac{F\cdot S}{q}=\frac{0.037\cdot 0.24}{35.9 \cdot 10^{-6}}=247.354 B"